# Sound Unblending: Exploring Sound Manipulations for Accessible Mixed-Reality Awareness

### v1

@@ 6. Example Applications with Sound Unblending
A prototype of an accessible online meeting application and demonstrated waveforms showing how sounds are manipulated. (a) The voices of the three people are assigned to the left, front, and right spatial locations around the BVI user, same as (b) the layout of the attendee panel. (c) When using a screen reader to navigate, the meeting volume is decreased to accentuate the audio feedback of the screen reader.
@@ 6.1. Accessible Online Meeting Application
In response to difficulties highlighted in section 3.2.1 about using screen readers during virtual meetings, we developed an audio-adaptive online meeting web application. This app resolves conflicts between screen reader and meeting audio using Prioritizer, seamlessly blending these sounds by adjusting their auditory characteristics. For example, it increases the screen reader’s volume when someone else speaks but allows users to prioritize meeting audio if it holds higher priority. To address the lack of spatial awareness common in virtual meetings, Spatializer arranges attendees’ voices across the left-right audio spectrum, corresponding to their positions on the attendee panel. This feature lets users easily map the voice’s location to the attendee’s location on the interface for further interaction (e.g., sending private messages).
@@ 6.2. Content-Aware MR Image Exploration
Mixed-reality images are increasingly pervasive, such as in news or research articles, which entail rich visual information. However, the spatial information is not adapted from the visual to the audio domain. In response, our image touch exploration system built on (Lee et al., 2022a) spatializes the screen reader’s voice feedback by mapping the touch location on the image to the sound feedback’s spatial location (Figure 10) for enhancing spatial understanding. Moreover, MR images, which combine digital and physical visuals easily distinguished visually but inaccessible to BVI people. To address this, during image exploration, Ambience Builder renders acoustic transparency for real-world (RW) content and occlusive opacity for virtual reality (VR) content, such as crowd noises in the office for RW content and bird chirping for VR content (Figure 10), which signals the reality and virtuality of content without additional text descriptions. Moreover, Stylizer alters the voice of description to a cartoonish tone to match the digital world’s style and enhance immersion.
@@
Using ImageExplorer (Lee et al., 2022a) with sound manipulations. The example image was provided by Wang et al. (Wang et al., 2020), which includes a combination of real and virtual content. The voice corresponding to the touch locations is spatialized from left to right. The voice of VR content is read with a cartoon-ish voice to fit the digital cartoon-style environment, while the voice of RW content remains as default. Noise cancellation mode is turned on when exploring VR content, and shifts to transparency mode when exploring RW content.
@@ 6.3. Context-Aware Outdoor Navigation
Navigation is indispensable to BVI people, and many commercial apps (sou, 2023; bli, 2023b) have been developed to facilitate BVI people’s independence and autonomy. However, occasional events may distract users from navigation. We developed a mobile navigation application that analyzes both real-world and virtual sounds to identify opportune moments for delivering audio directions. When a certain RW event is happening and detected, the audio directions will be delayed by Prioritizer until the RW event ends (Figure 11). If a certain RW sound is detected but overshadowed by the audio directions, this app can provide post-hoc verbal descriptions to notify users of the detected sound.
@@
Illustration of our context-aware navigation app. (a) The crowd passing by the BVI person generates sounds that are detected by the app, and at the same time, (b) an audio direction is about to happen but will not be played due to its conflict with crowd sounds. (c) After the crowd leaves, the audio direction is played.
@@ 7. Discussion and Future Work
We have contributed the concept of Sound Unblending with six sound manipulators to increase MR sound awareness for BVI people, a user study to prove the effectiveness of sound manipulations on different MR scenarios, and three example applications to demonstrate its feasibility for practical use. In this section, we discuss our work’s limitations and potential improvements, implications for future MR soundscape design, and the generalizability of our findings to broader communities.
@@ 7.1. From Simulation to Practical Applications
In our work, we created Unity simulations for real-time manipulation of sound characteristics and developed three proof-of-concept applications to illustrate the generalizability and practicality of Sound Unblending. We encountered several challenges; for instance, the audio streams in most existing platforms were at the operation system level, which made them hard to access for application-level purposes. Furthermore, implementing these sound manipulations in the real world necessitates several requirements, such as supporting the recognition of diverse categories of sounds, extracting sounds of interest from multiple overlapping ones, and rendering the extracted sounds in their original or higher quality. And all of these components should be achieved in real time to ensure a seamless user experience, which is still challenging in sound research. These are why our three example applications mostly revolve around manipulating virtual audio.
In recent years, researchers have attempted to tackle these challenges. For instance, Ubicoustics (Laput et al., 2018) supports the recognition of many sound activities through commercially available microphones, Jain et al. (Jain et al., 2022) developed ProtoSound for people to customize their sound recognition model, and Veluri et al. (Veluri et al., 2022) approached real-time target sound extraction. Also, to increase the accuracy of sound identification or extraction, we could approach the users’ sound context with other sensing modalities. For example, in the RW-Focused scenario, one’s smartwatch could provide gyroscope data to detect white cane taps. Additionally, everyday objects often produce sounds linked with visual elements, detectable through cameras, such as playing an instrument (Tian et al., 2021). These approaches can provide a preliminary context for the targeted sound identification or extraction. Despite a long way ahead in real-time manipulation of overarching sounds, recent research endeavors have illuminated promises to apply Sound Unblending to practical applications.
@@ 7.2. Towards Sound-Aware Description Manipulations
As the introduction outlines, incorporating visual descriptions is another essential auditory element to facilitate MR experiences for BVI people. Unlike diegetic sound effects (e.g., knocking, drilling, dog barking.), descriptions provide specific information and semantic meanings crucial for BVI users. In section 5.7.5, participants suggested making virtual broadcasts or voice notes discernible amidst other sounds only if the content is deemed critical, as a strategy to mitigate information overload. It is thus promising to adjust the description content or provide opportune sentence breaks to prevent it from overlapping with other MR auditory elements. For instance, the system can offer comprehensive descriptions during quieter moments, while providing succinct yet informative descriptions when time is limited, to create a harmonized user experience. This is similar to creating audio descriptions with dynamic time constraints in a video, such as Rescribe (Pavel et al., 2020) to shorten the descriptions by removing less important words. Overall, sound manipulation in MR environments should extend beyond adjusting sound characteristics, temporal, and spatial aspects. It should also consider the content and audio context, making it more intelligent, adaptive, and user-centric.
@@ 7.3. Customizable Sound Manipulations
Though each manipulator in our Unity implementation can manipulate sounds in real-time (e.g., overlap of several sounds.) to provide a seamless MR soundscape, we did not consider the varied preferences of our participants. As described in section 5.7.5, participants proposed several ideas on manipulating sounds based on their current context and content of audio information (e.g., sounds, text descriptions.), which also echos with our formative results in section 3.2.5 that BVI people desired to augment or customize the sound library of existing applications. It is thus promising to integrate methods that gauge the importance of audio information, such as content analysis or sentiment analysis. Based on the gauged importance, users could personalize sound representation (e.g., earcons, descriptions, diegetic sounds.) and presentation (e.g., high/low volume, sound locations, etc.) to minimize distraction. Additionally, participant preferences for sound manipulation, influenced by personal experiences and memories, highlight the need for adaptable soundscapes; for instance, the sound of the sliding door in the VR-Focused scenario made B1 recall her bad memory of consuming videos about insects, and she wanted to disable it selectively. Research on visual descriptions also emphasized individual customization over a one-size-fits-all approach (Kreiss et al., 2022; Stangl et al., 2021; Lopez et al., 2018, 2022). Echoing this notion, it is, therefore, worth exploring customizable sound properties and how to design end-user interfaces to enable user customization.
@@ 7.4. Generalizing Results to Broader Groups with Different Sensory Modalities
Furthermore, the simulations in our main study may not fully capture the multisensory experience that BVI individuals rely on, such as a combination of haptic feedback, smell, and other contextual cues, to enhance their real-world awareness. That was also why several participants in the RW-Focused scenario emphasized the importance of haptic feedback from the white cane besides its auditory feedback. Nonetheless, while other sensory feedback can augment and address the limitations of sound, they are all of significance and cannot replace one another. Also, none of our participants consumed visual feedback despite some having residual visual ability. Instead, they used hearing only to perform the tasks. Future work could investigate how people strategize their sensory modalities for balancing cognitive load and performance in mixed reality, which may enable manipulators across different modalities (e.g., split conflict audio information into visual and audio ones, and vice versa) to create more sensory-adaptive and accessible mixed reality experiences.
@@ 8. Conclusion
We have presented the concept of Sound Unblending to make MR sound awareness accessible for BVI people, through six sound manipulators derived from our content analysis on BVI forums, including Ambience Builder, Feature Shifter, Earcon Generator, Prioritizer, Spatializer, and Stylizer. We instantiated the six sound manipulators and three simulated scenarios across the Reality-Virtuality continuum in Unity. We then conducted a user study with eighteen BVI people and found empirical evidence that the six sound manipulations significantly enhanced users’ sound awareness and reduced cognitive load. We also found varied preferences and comments across participants on manipulating sounds, which spurred several discussions and promises of future work. Finally, we implemented three proof-of-concept applications to demonstrate the generalizability and practicality of Sound Unblending, including an accessible online meeting app, an immersive image understanding system, and a context-aware navigation app.

### v2

@@ 6. Example Applications with SoundShift
[…]
Figure 9 Figure 9a Joshua is a person who attends a virtual meeting with James, Bob, and Alice. This figure conceptually showing the the spatial audio enabling the atmosphere of physical meeting. Figure 9b The interface of the web application. There are attendee panels on the interface and some basic widgets. Figure 9c Waveforms show that the volume of screen reader will be louder than meeting when they are in conflict
@@ 6.1. Accessible Online Meeting Application
In response to difficulties highlighted in section 3.2.1 about using screen readers during virtual meetings, we developed an audio-adaptive online meeting web application. This app resolves conflicts between screen reader and meeting audio using Time Shift, seamlessly blending these sounds by adjusting their auditory characteristics. For example, it increases the screen reader’s volume when someone else speaks but allows users to prioritize meeting audio if it holds higher priority. To address the lack of spatial awareness common in virtual meetings, Position Shift arranges attendees’ voices across the left-right audio spectrum, corresponding to their positions on the attendee panel. This feature lets users easily map the voice’s location to the attendee’s location on the interface for further interaction (e.g., sending private messages).
@@ 6.2. Content-Aware MR Image Exploration
Mixed-reality images are increasingly pervasive, such as in news or research articles, which entail rich visual information. However, the spatial information is not adapted from the visual to the audio domain. In response, our image touch exploration system built on (Lee et al., 2022a) spatializes the screen reader’s voice feedback by mapping the touch location on the image to the sound feedback’s spatial location (Figure 10) for enhancing spatial understanding. Moreover, MR images, which combine digital and physical visuals easily distinguished visually but inaccessible to BVI people. To address this, during image exploration, Transparency Shift renders acoustic transparency for real-world (RW) content and occlusive opacity for virtual reality (VR) content, such as crowd noises in the office for RW content and bird chirping for VR content (Figure 10), which signals the reality and virtuality of content without additional text descriptions. Moreover, Style Shift alters the voice of description to a cartoonish tone to match the digital world’s style and enhance immersion.
@@
[…]
Figure 10 A smartphone showing a mixed reality image. The image contains several physical people wearing the head-mounted display. There are a open office space on the right of the image. Otherwise, people are seen as in the virtual park, which has digital cube-based trees and sky view.
@@ 6.3. Context-Aware Outdoor Navigation
Navigation is indispensable to BVI people, and many commercial apps (sou, 2023; bli, 2023b) have been developed to facilitate BVI people’s independence and autonomy. However, occasional events may distract users from navigation. We developed a mobile navigation application that analyzes both real-world and virtual sounds to identify opportune moments for delivering audio directions. When a certain RW event is happening and detected, the audio directions will be delayed by Time Shift until the RW event ends (Figure 11). If a certain RW sound is detected but overshadowed by the audio directions, this app can provide post-hoc verbal descriptions to notify users of the detected sound.
@@
[…]
Figure 11a A blind person is using a smartphone, and two people are passing by and talking to each other. There is text “Speech” showing on the top right corner of the phone screen, suggesting that the speech of the two people is detected. Figure 11b The phone screen shows the audio direction “Go right” in the top right corner with the red background, suggesting it is still pending Figure 11c After the two people leave, the audio direction “Go right” becomes green, suggesting that it is played to the user.
@@ 7. Discussion and Future Work
We have contributed the concept of SoundShift with six sound manipulators to increase MR sound awareness for BVI people, a user study to prove the effectiveness of sound manipulations on different MR scenarios, and three example applications to demonstrate its feasibility for practical use. In this section, we discuss our work’s limitations and potential improvements, implications for future MR soundscape design, and the generalizability of our findings to broader communities.
@@ 7.1. From Simulation to Practical Applications
In our work, we created Unity simulations for real-time manipulation of sound characteristics and developed three proof-of-concept applications to illustrate the generalizability and practicality of SoundShift. We encountered several challenges; for instance, the audio streams in most existing platforms were at the operation system level, which made them hard to access for application-level purposes. Furthermore, implementing these sound manipulations in the real world necessitates several requirements, such as supporting the recognition of diverse categories of sounds, extracting sounds of interest from multiple overlapping ones, and rendering the extracted sounds in their original or higher quality. And all of these components should be achieved in real time to ensure a seamless user experience, which is still challenging in sound research. These are why our three example applications mostly revolve around manipulating virtual audio.
In recent years, researchers have attempted to tackle these challenges. For instance, Ubicoustics (Laput et al., 2018) supports the recognition of many sound activities through commercially available microphones, Jain et al. (Jain et al., 2022) developed ProtoSound for people to customize their sound recognition model, and Veluri et al. (Veluri et al., 2023a) approached real-time target sound extraction. Also, to increase the accuracy of sound identification or extraction, we could approach the users’ sound context with other sensing modalities. For example, in the RW-Focused scenario, one’s smartwatch could provide gyroscope data to detect white cane taps. Additionally, everyday objects often produce sounds linked with visual elements, detectable through cameras, such as playing an instrument (Tian et al., 2021). These approaches can provide a preliminary context for the targeted sound identification or extraction. Despite a long way ahead in real-time manipulation of overarching sounds, recent research endeavors have illuminated promises to apply SoundShift to practical applications.
@@ 7.2. Towards Sound-Aware Description Manipulations
As the introduction outlines, incorporating visual descriptions is another essential auditory element to facilitate MR experiences for BVI people. Unlike diegetic sound effects (e.g., knocking, drilling, dog barking), descriptions provide specific information and semantic meanings crucial for BVI users. In section 5.7.5, participants suggested making virtual broadcasts or voice notes discernible amidst other sounds only if the content is deemed critical, as a strategy to mitigate information overload. It is thus promising to adjust the description content or provide opportune sentence breaks to prevent it from overlapping with other MR auditory elements. For instance, the system can offer comprehensive descriptions during quieter moments, while providing succinct yet informative descriptions when time is limited, to create a harmonized user experience. This is similar to creating audio descriptions with dynamic time constraints in a video, such as Rescribe (Pavel et al., 2020) to shorten the descriptions by removing less important words. Overall, sound manipulation in MR environments should extend beyond adjusting sound characteristics, temporal, and spatial aspects. It should also consider the content and audio context, making it more intelligent, adaptive, and user-centric.
@@ 7.3. Customizable Sound Manipulations
Though each manipulator in our Unity implementation can manipulate sounds in real-time (e.g., overlap of several sounds) to provide a seamless MR soundscape, we did not consider the varied preferences of our participants. As described in section 5.7.5, participants proposed several ideas on manipulating sounds based on their current context and content of audio information (e.g., sounds, text descriptions), which also echos with our formative results in section 3.2.5 that BVI people desired to augment or customize the sound library of existing applications. It is thus promising to integrate methods that gauge the importance of audio information, such as content analysis or sentiment analysis. Based on the gauged importance, users could personalize sound representation (e.g., earcons, descriptions, diegetic sounds) and presentation (e.g., high/low volume, sound locations) to minimize distraction. Additionally, participant preferences for sound manipulation, influenced by personal experiences and memories, highlight the need for adaptable soundscapes; for instance, the sound of the sliding door in the VR-Focused scenario made B1 recall her bad memory of consuming videos about insects, and she wanted to disable it selectively. Research on visual descriptions also emphasized individual customization over a one-size-fits-all approach (Kreiss et al., 2022; Stangl et al., 2021; Lopez et al., 2018, 2022). Echoing this notion, it is, therefore, worth exploring customizable sound properties and how to design end-user interfaces to enable user customization.
@@ 7.4. Generalizing Results to Broader Groups with Different Sensory Modalities
Furthermore, the simulations in our main study may not fully capture the multisensory experience that BVI individuals rely on, such as a combination of haptic feedback, smell, and other contextual cues, to enhance their real-world awareness. That was also why several participants in the RW-Focused scenario emphasized the importance of haptic feedback from the white cane besides its auditory feedback. Nonetheless, while other sensory feedback can augment and address the limitations of sound, they are all of significance and cannot replace one another. Also, none of our participants consumed visual feedback despite some having residual visual ability. Instead, they used hearing only to perform the tasks. Future work could investigate how people strategize their sensory modalities for balancing cognitive load and performance in mixed reality, which may enable manipulators across modalities (e.g., split conflicting audio information into visual and audio ones) to create more sensory-adaptive and accessible mixed reality experiences.
@@ 8. Conclusion
We have presented the concept of SoundShift to make MR sound awareness accessible for BVI people, through six sound manipulators derived from our content analysis on BVI forums, including Transparency Shift, Envelope Shift, Position Shift, Style Shift, Time Shift, and Sound Append. We instantiated the six sound manipulators and three simulated scenarios across the Reality-Virtuality continuum in Unity. We then conducted a user study with eighteen BVI people and found empirical evidence that the six sound manipulations significantly enhanced users’ sound awareness and reduced cognitive load. We also found varied preferences and comments across participants on manipulating sounds, which spurred several discussions and promises of future work. Finally, we implemented three proof-of-concept applications to demonstrate the generalizability and practicality of SoundShift, including an accessible online meeting app, an immersive image understanding system, and a context-aware navigation app.
Acknowledgements.
We thank our anonymous reviewers for their suggestions and all the participants in our study.